OpenStudy (anonymous):

What is the vertex form of the equation y=x^2-6x+8

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

No :/

jimthompson5910 (jim_thompson5910):

well first off, you start with this formula x = -b/(2a) and this will give you the axis of symmetry

jimthompson5910 (jim_thompson5910):

in this case, a = 1, b = -6 so that means x = -b/(2a) x = -(-6)/(2*1) x = ???

OpenStudy (anonymous):

x = -3 Not sure. I need a calculator

jimthompson5910 (jim_thompson5910):

close, but the two negatives cancel to make a positive

jimthompson5910 (jim_thompson5910):

so x = 3

jimthompson5910 (jim_thompson5910):

that's the axis of symmetry

jimthompson5910 (jim_thompson5910):

aka...the x coordinate of the vertex

jimthompson5910 (jim_thompson5910):

plug that x value into the equation to find the y coordinate of the vertex

OpenStudy (anonymous):

y=3^2-6(3)+8

OpenStudy (anonymous):

to find the y????

jimthompson5910 (jim_thompson5910):

yep, keep going

OpenStudy (anonymous):

y=9-18+8 y=9-26 ?

jimthompson5910 (jim_thompson5910):

-18+8 is NOT -26

OpenStudy (anonymous):

Well I kept the sign after 9. So it would be +?

jimthompson5910 (jim_thompson5910):

-18+8 is -10 think of it as: you have 18 IOU's, then you pay off 8 of them...so you're left with 10 IOUs

OpenStudy (anonymous):

-10

jimthompson5910 (jim_thompson5910):

9-18+8 9-10 = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

so the y coordinate is -1

jimthompson5910 (jim_thompson5910):

so what is the final answer?

OpenStudy (anonymous):

So I plug all that into the vertex form?

jimthompson5910 (jim_thompson5910):

the y coordinate of the vertex I meant to say

jimthompson5910 (jim_thompson5910):

yes you do

jimthompson5910 (jim_thompson5910):

a = 1 (given) h = 3 (x coordinate of vertex) k = -1 (y coordinate of vertex)